OpenStudy (anonymous):

For the simple harmonic motion equation d=5sin((π/4)t), what is the frequency?

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

d=Asin(ωt) ω=2πf, π/4=2πf, πf=π/8, f=1/8